Write the six steps in hypothesis testing in each of the following studies. Take note, the p-value is already given.


The average family size in the Philippines was reported as 4.25. A random sample of families in a particular street resulted in the following family sizes:


5, 6, 7, 4, 3, 8, 2 and 9


\At 0.10, is the average family size more than the national average?


(p-value = 0.0961)


1
Expert's answer
2022-06-03T13:04:58-0400

Sample mean 


"\bar{x}=\dfrac{1}{8}(5+6+7+4+3+8+2+9)=5.5"


Sample variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=6""s=\sqrt{6}\approx2.45"


The following null and alternative hypotheses need to be tested:

"H_0:\mu\le4.25"

"H_1:\mu>4.25"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=7" and the critical value for a right-tailed test is "t_c =1.4149236."

The rejection region for this right-tailed test is "R = \{t:t>1.414923\}."


The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{5.5-4.25}{2.45/\sqrt{8}}=1.443075"


Since it is observed that "t=1.443075>1.414923=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=7" degrees of freedom, "t=1.443075" is "p=0.0961," and since "p= 0.096107<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 4.25, at the "\alpha = 0.10" significance level.